OpenStudy (satsuki):

x^2-9y^2+2x-54y-89=0 Find the center, transverse axis, foci and asymtopes of the hyperbola

OpenStudy (anonymous):

complete the square first

OpenStudy (anonymous):

I assume you know how to do it. It turns out the equation is equivalennt to (x+1)^2-9 (y+3)^2-9 = 0 and so, [(x+1)^2]/3^2 - [(y+3)^2]/1^2 = 1

OpenStudy (anonymous):

yes? no?

OpenStudy (satsuki):

I did (x^2x+1)-6(y^2+6y+9)=89+1-36 to complete the square

OpenStudy (satsuki):

Where did you get the -9 from?